# arXiv:2603.01692v1[cs.LG] 02 Mar 2026

### v1

@@ C.3.1 Candidate Hypothesis Construction
Inspired by statistical physics (Isihara, 2013), we introduce a probabilistic interaction mechanism that enables controlled information exchange between traces. For each hypothesis $\eta_{c}$ generated by the current trace, we compute an interaction potential with all hypotheses $\eta_{j}$ in the shared memory $\mathcal{M}$:
$U_{cj}=\alpha\cdot S_{cj}\cdot e^{-\gamma L}+\beta\cdot\tanh(\Delta_{cj})\in[-2,2],$
@@
$p_{cj}=\frac{\exp(U_{cj})}{\sum_{k}\exp(U_{ck})},\quad\eta_{\text{sim}}\sim\text{Categorical}(p_{cj}).$
@@ C.4 Algorithm for Gome Main Loop
1:Input: task specification $\mathcal{T}$, dataset $\mathcal{D}$, number of traces $N$, time budget $B$
@@ D.1 Preventing Overfitting via Hierarchical Validation
A critical limitation of gradient-free optimization is the reliance on validation scores as the sole proxy for solution quality. This often leads to overfitting or the acceptance of technically flawed solutions that happen to score well on a specific split.
@@
Case: Overfitting Detection on Stanford COVID Vaccine (Node 32) Task: Predict RNA degradation rates (MCRMSE, lower is better) Hypothesis $\eta_{t}$: Align training objective with evaluation by redefining loss weights: for reactivity use clipped inverse-variance weights $w=\mathrm{clip}(1/(\mathrm{err}^{2}+10^{-6}),0.1,10)$; for deg_Mg_pH10 and deg_Mg_50C use constant $w=1.0$. Replace hard SN_filter gating with soft factor (1.0 for SN=1, 0.5 for SN=0). Per-batch renormalize so each scored channel has equal total weight mass. Metrics: Previous (Parent 27): $h^{*(i)}_{\mathrm{val}}=0.6948$, $h^{*(i)}_{\mathrm{test}}=0.2688$ Current (Node 32): $h_{t,\mathrm{val}}=0.2945$ ($\downarrow 57.6\%$ better), $h_{t,\mathrm{test}}=0.6382$ ($\uparrow 137.5\%$ worse)   Code Changes: Kept PairAwareBiGRU architecture, features, split, inference identical to SOTA. Replaced training weights with channel-specific weighting; introduced per-batch channel renormalization; left validation metric as unweighted MCRMSE. Observations from $f_{t}$: Current MCRMSE is 0.2945 vs. SOTA 0.2559; performance degraded by +0.0385+0.0385. EDA indicates weak target–error association for reactivity (Pearson $\approx-0.042$, Spearman $\approx 0.280$), suggesting inverse-variance weighting may not benefit learning. Hypothesis Evaluation: REFUTED. The redesigned loss weighting and per-batch renormalization worsened validation MCRMSE. Likely causes: weak correlation between reactivity and its error; soft SN scaling introducing noisier samples; renormalization potentially overemphasizing harder channels. Decision: $d_{t}=0$ (REJECT) — Despite validation appearing to improve, the structured analysis identified that the “improvement” stemmed from metric misalignment rather than genuine generalization gains.
@@
This case demonstrates why gradient-free methods that rely solely on scalar ranking are fundamentally limited: they cannot distinguish genuine improvements from deceptive ones. A score-driven agent would have accepted Loop 32’s solution based on the 57.6% validation “improvement,” leading to catastrophic test degradation. Gome’s structured feedback, which analyzes code changes, training dynamics, and feature correlations, provides the semantic information necessary to detect such patterns. The 66.7% detection rate, while not perfect, represents a significant advantage over purely score-driven approaches that would accept all 9 overfitting cases (0% detection).
D.2 Comparison with MLE-STAR
MLE-STAR (Nam et al., 2025) represents a recent advancement in MLE agents, achieving 63.6% medal rate on MLE-Bench Lite through web retrieval and targeted code-block refinement. While MLE-STAR and Gome share a chain-based structure (as opposed to tree or graph topologies), they differ fundamentally in how feedback guides optimization. We analyze these differences across three dimensions.
@@ D.2 Comparison with MLE-STAR
D.3 Online Kaggle Validation: Store Sales Forecasting
@@ D.4 Error Analysis
Appendix E Code and Artifacts
We release our codebase and solution traces to facilitate reproducibility and future research. The repository includes the full implementation of Gome with all prompt templates, as well as a representative subset of 40 optimization trajectories (in JSON format) from GPT-5 experiments due to file size constraints.

### v2

@@ C.3.1 Candidate Hypothesis Construction
Inspired by statistical physics [isihara2013statistical], we introduce a probabilistic interaction mechanism that enables controlled information exchange between traces. For each hypothesis $\eta_{c}$ generated by the current trace, we compute an interaction potential with all hypotheses $\eta_{j}$ in the shared memory $\mathcal{M}$:
$U_{cj}=\alpha\cdot S_{cj}\cdot e^{-\gamma L}+\beta\cdot\tanh(\Delta_{cj})\in[-2,2],$ (14)
@@
$p_{cj}=\frac{\exp(U_{cj})}{\sum_{k}\exp(U_{ck})},\quad\eta_{\text{sim}}\sim\text{Categorical}(p_{cj}).$ (16)
@@ C.4 Algorithm for Gome Main Loop
1:Input: task specification $\mathcal{T}$, dataset $\mathcal{D}$, number of traces NN, time budget $B$
@@ D.1 Preventing Overfitting via Hierarchical Validation
A critical limitation of score-centric feedback utilization is the reliance on validation scores as the sole proxy for solution quality. This often leads to overfitting or the acceptance of technically flawed solutions that happen to score well on a specific split.
@@
Case: Overfitting Detection on Stanford COVID Vaccine (Node 32) Task: Predict RNA degradation rates (MCRMSE, lower is better)
Hypothesis $\eta_{t}$: Align training objective with evaluation by redefining loss weights: for reactivity use clipped inverse-variance weights $w=\mathrm{clip}(1/(\mathrm{err}^{2}+10^{-6}),0.1,10)$; for deg_Mg_pH10 and deg_Mg_50C use constant $w=1.0$. Replace hard SN_filter gating with soft factor (1.0 for SN=1, 0.5 for SN=0). Per-batch renormalize so each scored channel has equal total weight mass.
Metrics:
Previous (Parent 27): $h^{*(i)}_{\mathrm{val}}=0.6948$, $h^{*(i)}_{\mathrm{test}}=0.2688$ Current (Node 32): $h_{t,\mathrm{val}}=0.2945$ ($\downarrow 57.6\%$ better), $h_{t,\mathrm{test}}=0.6382$ ($\uparrow 137.5\%$ worse)   Code Changes: Kept PairAwareBiGRU architecture, features, split, inference identical to SOTA. Replaced training weights with channel-specific weighting; introduced per-batch channel renormalization; left validation metric as unweighted MCRMSE.
Observations from $f_{t}$: Current MCRMSE is 0.2945 vs. SOTA 0.2559; performance degraded by +0.0385+0.0385. EDA indicates weak target–error association for reactivity (Pearson $\approx-0.042$, Spearman $\approx 0.280$), suggesting inverse-variance weighting may not benefit learning.
Hypothesis Evaluation: REFUTED. The redesigned loss weighting and per-batch renormalization worsened validation MCRMSE. Likely causes: weak correlation between reactivity and its error; soft SN scaling introducing noisier samples; renormalization potentially overemphasizing harder channels.
Decision: $d_{t}=0$ (REJECT) — Despite validation appearing to improve, the structured analysis identified that the “improvement” stemmed from metric misalignment rather than genuine generalization gains.
@@
This case demonstrates why methods that use feedback primarily for scalar ranking are fundamentally limited: they cannot distinguish genuine improvements from deceptive ones. A score-driven agent would have accepted Loop 32’s solution based on the 57.6% validation “improvement,” leading to catastrophic test degradation. Gome’s structured feedback, which analyzes code changes, training dynamics, and feature correlations, provides the semantic information necessary to detect such patterns. The 66.7% detection rate, while not perfect, represents a significant advantage over purely score-driven approaches that would accept all 9 overfitting cases (0% detection).
D.2 Cross-Task Analysis of Deceptive Improvements
Our earlier analysis focused on a single representative task. This section extends the analysis across multiple MLE-Bench tasks to identify systematic patterns in deceptive improvements—iterations where validation metrics improve but test performance degrades.
Task-Type Vulnerability.
Examining GPT-5 optimization traces across the full MLE-Bench, we find that deceptive improvements are not uniformly distributed but concentrate in two task categories:
  1. 1.
    High-dimensional sparse-feature tasks. Tasks such as RNA sequence modeling (stanford-covid-vaccine), genomic prediction, and high-cardinality categorical problems are most vulnerable. The high dimensionality and feature sparsity create opportunities for complex loss reweighting or subtle feature engineering to exploit spurious correlations that appear genuine on limited validation splits but fail to transfer to the test distribution.
  2. 2.
    Small-sample time-series tasks. Tasks with limited temporal data are prone to temporal feature engineering that overfits to idiosyncratic patterns in the training window. Highly specific lag features, holiday indicators, or trend decompositions can capture noise rather than signal, producing validation gains that do not transfer across the temporal boundary to the test period.
Modification Taxonomy.
Across the identified deceptive improvements, three dominant modification patterns emerge:
  1. 1.
    Loss reweighting and objective misalignment. The agent modifies the training loss (e.g., channel-specific weighting, sample reweighting, focal loss variants) in ways that align well with the validation metric on the current split but introduce systematic bias. This category is the most reliably detected by hierarchical validation, as the code changes concentrate in the training loop, making it relatively straightforward for the LLM validator to identify discrepancies between training objective and evaluation metric.
  2. 2.
    Aggressive feature engineering. The agent introduces highly specific features—interaction terms between rare categories, narrow time windows, task-specific hand-crafted indicators—that capture noise in the training/validation split but encode distributional artifacts rather than causal patterns. These are moderately detectable, as the validator can flag suspiciously specific features by analyzing their construction logic.
  3. 3.
    Simplicity Bias. After repeated failed iterations, the agent sometimes resorts to surface-level heuristics or hard-coded shortcut rules: threshold-based classifications mapping input ranges directly to output labels, median imputation strategies that “hack” the evaluation metric without learning the underlying data logic, or constant-prediction fallbacks exploiting metric asymmetries. This pattern aligns with the “shortcut solvability” phenomenon [nie2026dsgym], where agents prioritize superficial metric gains over genuine reasoning. Simplicity Bias is the hardest category to detect, because shortcut solutions often produce clean, syntactically reasonable code—the validator must reason about the semantic adequacy of the approach relative to the task complexity, which requires deeper understanding.
D.3 Comparison with MLE-STAR
MLE-STAR [nam2025mle] represents a recent advancement in MLE agents, achieving 63.6% medal rate on MLE-Bench Lite through web retrieval and targeted code-block refinement. While MLE-STAR and Gome share a chain-based structure (as opposed to tree or graph topologies), they differ fundamentally in how feedback guides optimization. We analyze these differences across three dimensions.
@@ D.3 Comparison with MLE-STAR
D.4 Online Kaggle Validation: Store Sales Forecasting
@@ D.5 Error Analysis
Appendix E Theoretical Motivation for the Smoothness Assumption
Gome’s gradient-based paradigm assumes that the MLE optimization landscape is locally smooth enough to support directed, incremental refinement. This section summarizes the theoretical rationale and empirical evidence for this assumption.
Loss-Surface Connectivity in Neural Networks.
A growing line of optimization research indicates that neural-network loss surfaces are substantially more benign than worst-case non-convex intuition implies. garipov2018loss show that independently trained optima can often be connected by simple low-loss paths (“mode connectivity”) rather than separated by high-loss barriers. bottou2018optimization further argue that stochastic gradient methods are effective in overparameterized regimes because optimization trajectories frequently traverse locally smooth, connected regions. Taken together, these results imply that moving from a good solution to a better one often proceeds through a sequence of small improvements rather than discontinuous jumps.
From Parameter Space to Code Space.
Gome optimizes code rather than parameters directly. The smoothness intuition transfers to code space because Gome’s structured reasoning does not propose arbitrary syntactic edits; instead, it uses diagnostics (training dynamics, validation curves, error distributions, and feature-importance patterns) to target changes in the training procedure—for example, learning-rate schedules, feature engineering, loss design, or regularization. These behavior-level modifications induce controlled changes in the underlying optimization trajectory. In this sense, Gome’s “gradient” operates on the pipeline’s functional behavior, not raw syntax. This is a weaker requirement than global smoothness of the code-to-performance mapping: it only requires that reasoning-guided edits produce locally predictable performance changes.
Empirical Evidence.
Our experiments provide practical evidence for local smoothness:
  • High per-iteration success rate. Table 5 reports a 41.1% improvement rate, meaning that nearly half of optimization steps yield validated gains through incremental updates. If the landscape were dominated by discontinuities or isolated basins, this rate would be much lower.
  • Sustained convergence trajectories. Figure 3(b) shows smooth, steep descent for Gome on Frontier-tier models, consistent with gradient-style optimization in locally smooth regions. By contrast, MCTS exhibits more stepwise, plateau-heavy dynamics that are typical of search in rougher landscapes.
  • Kaggle case study. In the Store Sales deployment (Appendix D.4), performance improves progressively over 48 iterations (RMSLE 0.485 $\rightarrow$ 0.318), with degrading modifications filtered out by the gate mechanism. This monotonic trajectory is consistent with iterative optimization along a smooth local path.
Handling Non-Smooth Transitions.
Some transitions in MLE are intrinsically discontinuous (e.g., switching from linear models to neural networks, or from single models to ensembles). Gome mitigates this with two mechanisms:
  1. 1.
    Multi-trace forced diversification3.5) initializes $N$ traces with distinct architectural hypotheses (e.g., gradient boosting vs. neural networks vs. ensembles), providing broad coverage of separate regions in solution space. This is analogous to multi-start optimization in classical non-convex settings [bottou2018optimization].
  2. 2.
    Cross-trace sharing via the probabilistic interaction kernel (Appendix C.3) allows traces to exchange validated strategies. When one trace discovers a better region, others can adopt or adapt that hypothesis, enabling non-local transitions while preserving local refinement within each region.
Scope.
The smoothness assumption is not universal. Tasks with extremely sparse rewards or pathological loss landscapes may violate local smoothness. However, in competitive data-science settings, top solutions often share backbone algorithms and differ mainly in feature engineering, hyperparameter tuning, and ensembling. This structure suggests that the MLE landscape is sufficiently smooth for gradient-based optimization to be effective in practice. The scaling analysis in §5 provides further indirect support: the monotonic widening of Gome’s advantage with model capability is consistent with stronger reasoners producing higher-fidelity gradient signals that better exploit local smoothness.